# Optical Image-to-Image Translation Using Denoising Diffusion Models: Heterogeneous Change Detection as a Use Case

### v1

@@ Abstract
We introduce an innovative deep learning-based method that uses a denoising diffusion-based model to translate low-resolution images to high-resolution ones from different optical sensors while preserving the contents and avoiding undesired artifacts. The proposed method is trained and tested on a large and diverse data set of paired Sentinel-II and Planet Dove images. We show that it can solve serious image generation issues observed when the popular classifier-free guided Denoising Diffusion Implicit Model (DDIM) framework is used in the task of Image-to-Image Translation of multi-sensor optical remote sensing images and that it can generate large images with highly consistent patches, both in colors and in features. Moreover, we demonstrate how our method improves heterogeneous change detection results in two urban areas: Beirut, Lebanon, and Austin, USA. Our contributions are: i) a new training and testing algorithm based on denoising diffusion models for optical image translation; ii) a comprehensive image quality evaluation and ablation study; iii) a comparison with the classifier-free guided DDIM framework; and iv) change detection experiments on heterogeneous data.
@@ I Introduction
Image-to-image translation (I2I) is a technique that transforms an image from one domain to another while preserving its original content [Saharia2021, Dhariwal2021, I2ISeo2023, Park2019, I2IIsmael2023, I2ITasar2020, I2IZhang2023, I2ISokolov2023]. This is useful for various applications such as super-resolution, inpainting, image decompression, and domain adaptation. In the context of remote sensing, I2I is useful when images acquired from different sensors need to be compared. For example, a deep learning model that learns from paired images can translate a low-resolution optical image to a high-resolution optical image or vice versa. Thus, synthetic versions of high-resolution images from low-resolution ones can be created, which are useful when the original high-resolution images are not available or too costly to acquire. However, I2I in this context is challenging because of the differences in spatial and spectral characteristics, noise levels, and geometric distortions of the heterogeneous images. Therefore, an I2I model needs to learn how to preserve the content and structure of the source image while adapting to the style and resolution of the target domain.
We address the shortcomings of existing I2I methods [CDFang2023, I2IZhang2023, I2ISeo2023, CDLi2021, Xiao2023] for remote sensing imaging, which focus on patch-level translation and do not account for the variability of optical images due to multiple factors. Also, so far, no other work has proposed an algorithm that performs domain adaptation and super-resolution as a singular task using denoising diffusion models (DDM). A novel optical-to-optical translation approach is hereby propose. It leverages Denoising Diffusion Models (DDM) [Ho2020], which have superior image generation performance and avoid the problems of GANs and their variants [mansourifar_i3eaccess2022], such as mode-collapse and undesired hallucinations. We aim to achieve domain-robust and high-quality I2I translation using DDMs, which are used by several leading AI companies for image generation.
@@
The proposed DDM-based method differs from the standard framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022], as we have observed that the latter produces results that are highly inconsistent with neighboring patches, creating checkerboard-like images when patches are combined for a larger area. In contribution to this effect, the classifier-guided Denoising Diffusion Implicit Model (DDIM) [ho2022classifierfree, Song2020] seems to get confused by the high variability of colors between training patches when their features are very similar. This causes the generation of patches whose overall colorization is highly inaccurate. Our method tackle this problem by using color-standardization pre and post processing procedures carefully tailored to the optical-to-optical translation task.
@@
Moreover, due to its stochastic nature, denoising diffusion models produce different outputs for a given input [Ho2020]. Even when not adding noise in between reverse diffusion steps [Song2020], different choices of the initial random noise matrix result in different outputs for the same input. This characteristic, while extremely useful for tasks like inpainting and unconditional generation, can bring uncertainty about the consistency of generations in the context of I2I translation for remote sensing. In our experiments, we observed that the randomness brought by the initial noise matrix translates into the creation of a few but non-negligible low-quality generations caused by bad choices of initial noise conditions. These generations can be inconsistent with neighboring patches and with the features present in its correspondent input patch. To alleviate this, we propose a modification in the DDIM inference procedure that optimizes the choice of the initial random noise matrix.
@@
As experiments, we apply our DDM-based method to generate synthetic high-resolution images for eight regions of interest: Beirut, Austin, Tlaquepaque, and five other locations. We evaluate the quality of the images produced by the proposed model using Learned Perceptual Image Patch Similarity (LPIPS) [Zhang2018LPIPS], Fréchet Inception Distance (FID) [Heusel2017], and PSNR [Hore2010]. We find that our I2I method shows great improvement in image quality when using Sentinel-II as input, compared to the original Sentinel-II images and to synthetic images generated by deep regression-based models. To test the method in a practical application, we perform two heterogeneous change detection (HCD) experiments. HCD is a technique that compares images acquired from different sensors or different modes of image acquisition to identify changes in a specific area. Valuable for quick disaster damage assessment, it is, however, challenging to achieve accurate and consistent results due to the differences in image characteristics and quality [Chini2023, Touati2020, Zhang2022, WangCD2022, Lv2022, Wu2022]. The presented HCD tests show how our proposed I2I method can alleviate the effect of such differences. These experiments are performed on two regions: Beirut, Lebanon, and Austin, Texas, using a targetless change detection algorithm based on classical image processing tools. We compare the HCD results with and without the synthetic images. Our results demonstrate that the synthetic images generated by the proposed I2I method improve the HCD performance by greatly reducing false alarms and enhancing the change regions. Moreover, the synthetic images generated for this test are realistic, detailed, consistent, and with better contrast and clarity, compared to the original low-resolution images.
We summarize the main contributions of this work as:
@@
    The proposal of training and testing procedures based on denoising diffusion models that are able to translate optical images from a lower spatial resolution to a higher spatial resolution domain while maintaining high inter-patch and input-output consistency.
@@
    The presentation of image quality comparisons between the proposed method and other solutions, together with ablation experiments, over hundreds of square kilometers of imagery.
@@
    The execution of tests that show how one of the most ubiquitous frameworks for diffusion models, the classifier-free guided DDIM [ho2022classifierfree], might not be well suited to perform I2I translation of optical remote sensing images when trained with large and diverse data sets, and how our method is not affected by the same problems.
  4. 4.
    The presentation of heterogeneous change detection tests in Beirut, Lebanon, and Austin, USA, indicate that the proposed I2I translation method has the potential to improve heterogeneous change detection performance, compared to when no translation between domains is performed.
I-A Related Works
I-A1 Denoising Diffusion Models
Ho et al. [Ho2020] proposed the use of a denoising-based framework for image generation, the so-called Denoising Diffusion Probabilistic Model (DDPM). It showed incredible image generation quality and overcame most state-of-the-art methods available at the time. However, the inference process was highly inefficient since hundreds of model evaluations per image were necessary to reach optimal performance. The work Denoising Diffusion Implicit Models (DDIM) [Song2020] approached this issue by presenting a new interpretation of the reverse diffusion process. Differently from the work in [Ho2020], they treated reverse diffusion as a non-Markovian process by skipping multiple reverse diffusion steps. Furthermore, they experimented with eliminating the randomness of intermediate diffusion steps—which came from the addition of a random noise matrix to each reverse diffusion step result—and noticed that this enabled an even greater reduction of model evaluations during inference at the cost of sample variety. In practice, these combined modifications reduced the necessary number of model evaluations during inference by up to two orders of magnitude while maintaining high-generation quality.
I-A2 Image-to-Image Translation
Other works focused on remote sensing-related tasks. Ismael et al. [I2IIsmael2023] proposed a GAN-based I2I translation method with the goal of performing remote sensing semantic segmentation of optical images from unknown domains using a segmentation network trained with optical images from a single domain. The methods in [I2ITasar2020] and [I2ISokolov2023] proposed style-transferring domain adapting frameworks that generate extra training data matching the appearance of a particular same-resolution domain, to alleviate train/test data drift. The method in [CDLi2021] uses a cycle-consistent GAN to translate optical images into SAR images, to perform change detection. It showed promising change detection results. However, training and testing data are limited to just a few acquisitions. The method in [I2IZhang2023] is a thermodynamics-inspired translation network that obtained interesting results in SAR-to-Optical translation tasks with its unconventional design. The work in [I2ISeo2023] proposes a DDM-based Brownian Bridge method for SAR-to-Optical I2I translation. There, the authors suggest the use of a modified forward diffusion procedure that, instead of resulting in a pure-noise image, the target SAR image is iteratively transformed into the source optical image. They reason that this could help the model more easily understand the relationships across domains. Xiao et al. propose a diffusion-based super-resolution method that seeks to further explore prior information from the low-resolution input. When tested on a small-scale data set, it performed well.
Despite the advancements in image-to-image translation within the context of remote sensing, existing literature predominantly focuses on the generation of additional training data for direct use in posterior deep learning algorithms [I2ITasar2020, I2IIsmael2023, I2ISokolov2023], or in SAR to Optical/Optical to SAR translation [CDLi2021, I2IZhang2023, I2ISeo2023]. These methods, while highly valuable, do not address the challenge of translating images across disparate optical sensors with different spatial resolutions, particularly ensuring feature consistency between input and output images. This gap is significant as it pertains to the fidelity of the translated images, where the preservation of salient features without introducing non-existent elements is crucial. Our work pioneers in this direction by proposing a novel framework that not only translates images from varying-resolution optical sensors but also guarantees the consistency of features. This approach mitigates the common issue of model hallucinations, where the generation of artifacts or the alteration of key elements could lead to misinterpretations, especially in critical applications like environmental monitoring and urban planning. By maintaining strict adherence to feature consistency, our method ensures that the output images faithfully represent the true characteristics of the input.
@@ I-A2 Image-to-Image Translation
Figure 1: Satellite images from Beirut, Lebanon, of low and high resolution, captured at a very similar time interval, from Sentinel-II and Planet Dove sensors. Figures 1a and 1c are the pre and post-event images from Sentinel-II, respectively. Figures 1b and 1d are the pre and post-event images from Planet Dove, respectively.
@@ II Proposed Method
Our method tackles the challenging task of optical-to-optical image translation between a lower spatial resolution sensor domain and a higher spatial resolution sensor domain. While higher spatial resolution is the most noticeable difference between such domains, there are also many other subtle features that might be difficult to pinpoint. The method hereby described leverages denoising probabilistic diffusion models to not only virtually increase spatial resolution but also to match these less obvious but important differences between the domains. This requires a training dataset with pairs of co-registered low spatial resolution (LR) domain $A$ and high spatial resolution (HR) domain $B$ images, showing the same scenes taken with different sensors, to ensure a thorough domain adaptation.
The full method generates multi-patch images with highly consistent and color-accurate patches. These images are sharp and clear, showing significant visual resolution improvements compared to lower-resolution domain images while preserving the original content. Our experiments in Section V provide evidence of these improvements.
@@
10:  $\gamma\sim\text{Uniform}(\bm{\gamma})$
11:  $\tilde{\bm{y}}\leftarrow\sqrt{\gamma}\bm{y}_{0}+\sqrt{1-\gamma}\bm{\varepsilon}$Forward diffusion}
12:  $\hat{\bm{\varepsilon}}\leftarrow\bm{\theta}_{\tau}\left(\tilde{\bm{y}},\bm{x}\right)$
@@ II-A Training
Our training procedure diverges from conventional Denoising Diffusion Probabilistic Models (DDPM) by incorporating classifier-free guidance training [ho2022classifierfree]. This addition enforces alignment between input and output patches, thereby mitigating the risk of model hallucinations. Additionally, we employ color standardization, a pre-and a post-processing technique: Whitening and Coloring, respectively. These steps are critical as they prevent the diffusion network from being confounded by the high variability in patch coloration found within extensive training datasets, particularly when the patches have highly similar content. Without these measures, the network struggles to accurately predict the channel mean values for the output patches, as shown in Section V.
@@
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the same features. Thus, patches aligned in time are required for training to work. Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$, which are, respectively, the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference, and the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast regions that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch. Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix, as also defined in [ho2022classifierfree]. To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until the learning has converged.

### v2

# Patch-Consistent Optical Translation Across Sensors: Large-Scale Denoising Diffusion with Heterogeneous Change Detection as a Use Case

@@ Abstract
In the field of remote sensing, the challenge of comparing images captured by disparate sensors is a common obstacle. This requires image translation—converting imagery from one sensor domain to another while preserving the original content. Denoising Diffusion Implicit Models (DDIM) are potential state-of-the-art solutions for such domain translation due to their proven superiority in multiple image-to-image translation tasks in classic computer vision. However, these models struggle with large-scale multi-patch imagery, often focusing solely on small patches and resulting in inconsistencies across the full image. To overcome these limitations, we propose a novel method that leverages DDIM for effective optical image translation over large areas. Our approach is tailored to super-resolve large-scale low spatial resolution images into high-resolution equivalents from disparate optical sensors, ensuring uniformity across hundreds of patches. Extensive experiments with a dataset of paired Sentinel-II and Planet Dove images show that our approach provides precise domain adaptation and artifact reduction. Our technique preserves the image content while also improving radiometric (color) accuracy and feature representations. The outcome is a high-resolution large-scale image with consistent patches, vital for applications such as heterogeneous change detection (HCD). We present a unique training and testing algorithm rooted in DDIMs, a thorough image quality assessment, and a comparative study against the standard classifier-free guided DDIM framework and five other leading methods. The efficacy of our approach is further demonstrated by substantial enhancements in HCD tasks performed in the urban settings of Beirut, Lebanon, and Austin, USA.
@@ I Introduction
Image-to-image translation (I2I) is the task of transforming an image belonging to one data domain to another domain while preserving at least partially its original content. This is useful for various applications such as super-resolution, inpainting, image decompression, and domain adaptation [Saharia2021, Dhariwal2021, I2ISeo2023, Park2019, I2IIsmael2023, I2ITasar2020, I2IZhang2023, I2ISokolov2023]. In the context of remote sensing, I2I is useful when images acquired from different sensors need to be compared. For example, a deep learning model that learns from paired images can translate a low-resolution domain $A$ optical image to a high-resolution domain $B$ optical image or vice versa. Thus, synthetic versions of high-resolution images from low-resolution ones can be created, which are useful when the original high-resolution images are not available, too costly to acquire, or plagued by unfavorable weather conditions. However, I2I in this context is challenging because of the differences in spatial and spectral characteristics, noise levels, and geometric distortions that are particular to each optical sensor. Therefore, an I2I model needs to learn how to preserve the content and structure of the source image while adapting to the radiometric properties and spatial resolution of the target domain.
We address the limitations of existing Image-to-Image (I2I) methods [CDFang2023, I2IZhang2023, I2ISeo2023, CDLi2021, Xiao2023] in remote sensing imaging, which are primarily designed for patch-level translation and fail to consider the variability of optical images influenced by various factors. Notably, our work is the first to propose an algorithm that integrates large-scale domain adaptation and super-resolution into a single task using Denoising Diffusion Models (DDM) [Ho2020].
Our novel optical-to-optical translation approach leverages the strengths of DDMs. These models have demonstrated exceptional image generation capabilities, surpassing adversarial models without encountering their common training issues, such as mode-collapse [mansourifar_i3eaccess2022]. The method hereby proposed exclusively employs Denoising Diffusion Implicit Models (DDIM), which are advantageous due to their reduced required number of model evaluations when compared to Denoising Diffusion Probabilistic Models (DDPM) and due to their deterministic behavior in intermediate inference steps. This approach ensures efficient and consistent domain-robust large-scale I2I translation, setting our method apart from existing techniques.
@@
The proposed method differs from the standard classifier-free guided DDIM framework commonly used in traditional computer vision tasks [ho2022classifierfree, Saharia2021, Dhariwal2021, Song2020, Nichol2021, Wang2022], as, during inference, it seems to get confused by the high variability of colors between training patches even when their features are very similar. This causes the generation of patches with whose overall colorization (radiometry) is highly inaccurate. Our method tackles this problem by using color-standardization pre- and post-processing procedures carefully tailored to the optical-to-optical translation task.
@@
Moreover, due to its stochastic nature, denoising diffusion models produce different outputs for a given input [Ho2020]. Even when not adding noise in between reverse diffusion steps, as in DDIM [Song2020], different choices of the initial random noise matrix result in different outputs for the same input. This characteristic, while crucial for tasks like inpainting and unconditional generation, can bring uncertainty about the consistency of generations in the context of I2I for remote sensing. In our experiments, we observed that the randomness brought by the initial noise matrix translates into the creation of a few but non-negligible low-quality generations caused by sub-optimal choices of initial noise conditions, leading to inconsistent feature generations across neighboring patches. To alleviate this, we propose a modification in the DDIM inference procedure that optimizes the choice of the initial random noise matrix.
@@
As experiments, we apply our DDM-based method to generate synthetic high-resolution images for eight regions of interest: Beirut, Austin, Tlaquepaque, and five other locations. Our I2I method greatly improves the image quality when using Sentinel-II as input, compared to the original Sentinel-II images and to synthetic images generated by other state-of-the-art models. To test the method in a practical application, we perform two heterogeneous change detection (HCD) experiments. HCD is a technique that compares images acquired from different sensors or different modes of image acquisition to identify changes in a specific area. Valuable for quick disaster damage assessment, it is, however, challenging to achieve accurate and consistent results due to the differences in image characteristics and quality [Chini2023, Touati2020, Zhang2022, WangCD2022, Lv2022, Wu2022]. The presented HCD tests show how our proposed I2I method can alleviate the effect of such differences. These experiments are performed on two regions: Beirut, Lebanon, and Austin, Texas, using a targetless change detection algorithm based on classical image processing tools. We compare HCD results obtained with and without synthetic images. The results demonstrate that the synthetic images generated by the proposed I2I method improve the HCD performance by greatly reducing false alarms and enhancing the change regions. Moreover, the synthetic images generated for this test are realistic, detailed, consistent, and with better contrast and clarity compared to the original low-resolution images.
Our contributions are summarized as follows:
@@
    We popose original training and testing procedures based on denoising diffusion models that are able to translate optical images from a low spatial resolution domain to a higher one over large areas while maintaining high inter-patch and input-output consistency.
@@
    The presentation of image quality comparisons between the proposed method and other cutting-edge solutions, together with ablation experiments, over hundreds of square kilometers of imagery, that highlights the contribution of our approach in contrast to the classical classifier-free guided DDIM one [ho2022classifierfree].
@@
    The presentation of heterogeneous change detection tests in Beirut, Lebanon, and Austin, USA, indicating the contribution of our proposed I2I translation method to heterogeneous change detection performance.
@@
Figure 1: Surface reflectance imagery from Beirut, Lebanon, of low and high spatial resolutions, captured at similar time intervals, from Sentinel-II (10m) and Planet Dove (3m) sensors. Figures 1a and 1c are the pre and post-event images from Sentinel-II, respectively. Figures 1b and 1d are the pre and post-event images from Planet Dove, respectively.
@@ II Proposed Method
Our method tackles the challenging task of large-scale patch-consistent optical-to-optical image translation from a lower spatial resolution sensor domain to a higher spatial resolution one. While spatial resolution is the most noticeable difference between such domains, there are also other disparities [CuiColor2020] such as radiometric fidelity and consistency, illumination conditions, and other factors such as sensor spectral sensitivity that increases furthermore the radiometric discrepancy between the two domains. The method hereby described leverages denoising probabilistic diffusion models (DDPMs) to not only virtually increase the spatial resolution but also to align subtle yet significant differences between the two domains. This requires a training dataset with pairs of co-registered low spatial resolution (LR) domain $A$ and high spatial resolution (HR) domain $B$ images, showing the same scenes taken with different sensors, to ensure proper domain adaptation.
The full method generates large multi-patch images with highly consistent and color-accurate patches, showing significant visual resolution improvement compared to lower-resolution domain images while preserving the original content (see Section IV). The model architecture and the associated hyperparameters used for our experiments are discussed in Appendix B. Details about the denoising diffusion working principles are presented in Appendix C.
@@
10:  $\gamma_{t}\sim\text{Uniform}(\bm{\gamma})$
11:  $\bm{y}_{t}\leftarrow\sqrt{\gamma_{t}}\bm{y}_{0}+\sqrt{1-\gamma_{t}}\bm{\varepsilon}$Forward diffusion}
12:  $\hat{\bm{\varepsilon}}\leftarrow\bm{\theta}_{\tau}\left(\bm{y}_{t},\bm{x}\right)$
@@ II-A Training
Our training procedure diverges from the conventional DDPM by incorporating classifier-free guidance training [ho2022classifierfree]. This addition enforces alignment between input and output patches. Additionally, we employ pre and post-processing techniques: Whitening and Coloring, respectively. These are critical as they prevent the diffusion network from being confounded by the high variability in patch coloration found within extensive training datasets, particularly when patches have highly similar content. Without these measures, the network fails to accurately predict the channel mean values for the output patches, as shown in Section IV. Details of the implementation of DDPM, DDIM, and Classifier-Free Guidance are respectively presented in Appendices C-A, C-B, and C-C.
Algorithm 1 defines the training step for a single pair of low spatial resolution patch from domain $A$, $\bm{I}_{\text{in}}^{\text{LR}}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and high resolution patch, $\bm{I}_{\text{out}}^{\text{HR}}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, from domain $B$. The goal is to train a model $\bm{\theta}$ that is able to convert $\bm{I}_{\text{in}}^{\text{LR, local}}$ into $\bm{I}_{\text{out}}^{\text{HR}}$. The training step is illustrated by Figure 2. The input and output images have the same pixel resolution but differ in spatial resolution, i.e., $\bm{I}_{\text{in}}^{\text{LR, local}}$ is up-scaled with cubic interpolation to have the pixel resolution of $\bm{I}_{\text{out}}^{\text{HR}}$, while covering the exact same area and the similar features. Thus, patches aligned in time are required for training to work.
Line 1 of the algorithm defines $\bm{I}_{\text{in}}^{\text{LR}}$ as the channel-wise stacking of $\bm{I}_{\text{in}}^{\text{LR,local}}$ and $\bm{I}_{\text{in}}^{\text{LR, global}}$. The matrix $\bm{I}_{\text{in}}^{\text{LR,local}}$ is defined as the region of interest (ROI) of the low-resolution image from domain $A$, which is the patch to be translated to domain $B$ in inference. $\bm{I}_{\text{in}}^{\text{LR, global}}$ represents the surroundings of the region of interest, down-sampled to have the same dimensions of $\bm{I}_{\text{in}}^{\text{LR,local}}$. $\bm{I}_{\text{in}}^{\text{LR, global}}$ is defined to contain four times the visible area of $\bm{I}_{\text{in}}^{\text{LR,local}}$, i.e., the area that is contained in $\bm{I}_{\text{in}}^{\text{LR,local}}$ plus three additional equally sized regions in its vicinity, thus depicting north, east and southeast patches that are connected to $\bm{I}_{\text{in}}^{\text{LR,local}}$. This is done to give more context to the model of what exactly lies around the input patch.
@@
Line 2 samples a random variable $u$ from a Bernoulli distribution of probability $p_{\text{uncond}}$. If it turns out to be $1$, the model behaves as an unconditional diffusion model for the current training step (Line 4), as in [ho2022classifierfree]. Otherwise, it behaves as a conditional diffusion model. ${\O}$ is the null-condition matrix (like in [ho2022classifierfree]). To ensure that ${\O}$ is correctly interpreted by the model, we set it as a matrix whose all elements are equal to $-2$, which is outside the range of values of input pixels $[-1,1]$. Thus, the model does not interpret the null condition as a regular input image and vice-versa. Lines 6 and 8 define the conditioning input from domain $A$, $\bm{x}\in\mathds{R}^{2n_{\text{ch}}\times h\times w}$, and the domain $B$ ground truth $\bm{y}_{0}\in\mathds{R}^{n_{\text{ch}}\times h\times w}$, respectively. They are obtained by applying the function $\mathbf{W}$, defined in Algorithm 2, to $\bm{I}_{\text{in}}^{\text{LR}}$ and $\bm{I}_{\text{out}}^{\text{HR}}$. This function extracts out the overall color information from an input image, i.e., it zeroes the mean of each channel. A normalization between $[-1,1]$ follows. $\mathbf{W}$ also returns the previous means, minimum, and maximum values extracted from the input, which can be used for reverting the whitening process during inference. However, during training, these values are not used. Lines 9 to 13 of Algorithm 1 are as in the original conditional DDPM training procedure [Saharia2021]. This process is repeated for all existing training patches until convergence.